OpenStudy (anonymous):

can someone actually give me a better answer with this question? 2.3x10^(4 + 4.4x10^(3

OpenStudy (anonymous):

is this \[2.3\times 10^4+4.4\times 10^3\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then \[2.3\times 10^4=23000\] and \[4.4\times 10^3=4400\] and when you add them you get \[23000+4400=27400\]

OpenStudy (anonymous):

is there a way to write that in scientific notation?

OpenStudy (amistre64):

.... i dont think so. scientific notation can only be used on numbers that are not this :/ otherwise, yes

OpenStudy (anonymous):

yes you can write \[2.74\times 10^4\]

OpenStudy (anonymous):

Okay thank you! i think thats it....

OpenStudy (anonymous):

i like when the calculator gets stuck i scientific notation mode so 23 comes out as \[2.3\times 10^1\]

OpenStudy (anonymous):

so the answer would be with a 1 as the exponent or a 4 as an expoonent